OpenStudy (anonymous):

the equation x^3 - 6x^2 + 11x - 3 = 0 has a solution near x = 0. Use Newton's Method to find an approx value (out of 3 decimal places) for this solution.

myininaya (myininaya):

It says the solution is near x=0. let's use an approximation for the solution. We make the initial guess 1/2.

myininaya (myininaya):

f'(x)=3x^2-12x+11

myininaya (myininaya):

x_1=x_0-f(x_0)/f'(x_0)=1/2-f(1/2)/f'(1/2)=...

OpenStudy (anonymous):

is it x_o - f(x_o) ALL OVER f'(x_o)? Or just f(x_o) / f'(x_o)

myininaya (myininaya):

you don't know the forumla? is that all you need help on? ok well its x-f/f'

myininaya (myininaya):

the x is not being divided by f' the f' is just under the f

OpenStudy (anonymous):

So after I plug in the 1/2, is that my answer? Or do I continue to get smaller like 1/4, 1/8, etc?

OpenStudy (anonymous):

Also, do I do negatives?

myininaya (myininaya):

no it says to do it til you get within 3 decimal points so when you do this you do it over and over until you see the same 3 decimal points repeating

myininaya (myininaya):

you can do negatives

OpenStudy (anonymous):

Thanks! Thank you so much. Have a great night!